OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (7, -2) and has a slope of -3 1. y - 2 = -3(x + 7) 2. y + 7 = -3(x - 2) 3. y + 2 = -3(x - 7) 4. y - 7 = -3(x + 2)

OpenStudy (anonymous):

number 3. Use y-y1=m(x-x1) y-(-2)=-3(x-7)

OpenStudy (anonymous):

oh i see, thank you very much

OpenStudy (anonymous):

no problem :)